# Adaptive Contrastive Search: Uncertainty-Guided Decoding for Open-Ended Text Generation

### v1

@@ Abstract
Decoding from the output distributions of large language models to produce high-quality text is a complex challenge in language modeling. Various approaches, such as beam search, sampling with temperature, $k-$sampling, nucleus $p-$sampling, typical decoding, contrastive decoding, and contrastive search, have been proposed to address this problem, aiming to improve coherence, diversity, as well as resemblance to human-generated text. In this study, we introduce adaptive contrastive search, a novel decoding strategy extending contrastive search by incorporating an adaptive degeneration penalty, guided by the estimated uncertainty of the model at each generation step. This strategy is designed to enhance both the creativity and diversity of the language modeling process while at the same time producing coherent and high-quality generated text output. Our findings indicate performance enhancement in both aspects, across different model architectures and datasets, underscoring the effectiveness of our method in text generation tasks. Our code base, datasets, and models are publicly available.
@@ 1 Introduction
The Transformer (Vaswani et al., 2017) plays a key role in various generative natural language processing (NLP) tasks such as generating stories, completing contextual text, and dialogue systems. However, the conventional method of training these models using maximum likelihood estimation (MLE) and decoding to the most probable sequence often results in substantial shortcomings. This can lead to repetitive and uncreative text generation, a problem known as degeneration.
Figure 1: Visualization of the adaptive contrastive search process: A three-step procedure utilizing entropy as a proxy for model uncertainty to guide the automated adaptation of the contrastive search parameters.
@@
We address this limitation by proposing adaptive contrastive search (ACS), an adaptive approach that automatically adjusts the hyperparameters of conventional CS. This method evaluates the model’s uncertainty at each generation step and adjusts the weighting of both components, without the need for manual intervention. The experimental outcomes demonstrate the effectiveness of our method, as it performs well in the task of open-ended text generation across different architectures and datasets.
@@ 2 Related work
These approaches focus on choosing the text continuation with the highest probability according to the model’s probability distribution. Prominent examples are beam search and greedy search. Recently, studies by Shao et al. (2017); Vijayakumar et al. (2018); Paulus et al. (2017); Klein et al. (2017) have demonstrated that solely maximizing the output probability frequently leads to degenerated or repetitive text sequences, problem that has been addressed by stochastic, sampling-based methods.
@@
Our study, however, focuses on the work of Su et al. (2022), where they introduce CS, where given the prompt text $\boldsymbol{x}_{<t}$, the selection of the output token $x_{t}$ follows:
@@
Empirical results suggest different values for $\alpha$ and $k$, depending on the task, the datasets, and the language of interest, respectively (Su et al., 2022; Su and Xu, 2022; Su and Collier, 2023).
An empirical study comparing CS and CD Su and Collier (2023) highlights the strengths and weaknesses of both approaches. The automatic evaluation results suggest that CD performs better on MAUVE Pillutla et al. (2021), while CS excels in diversity and coherence. Additionally, through extensive human evaluations, they demonstrate that human annotators universally prefer CS over CD by substantial margins. Given the contradictory results between MAUVE and human evaluations, their analysis reveals that balancing diversity and coherence metrics better correlates with human judgments.
@@ 3.2 Theoretical Motivation
Table 1: Automatic evaluation results: Numbers marked with * are obtained using the generated texts originally released by Li et al. (2023), CS results are taken from Su and Xu (2022). The best results are highlighted in bold.
@@ 4.1 Evaluation Metrics
Proposed by Su et al. (2022), coherence is estimated by the cosine similarity between the sentence embeddings of prompt $\mathrm{x}_{\text{prompt }}$ and generated continuation $\mathrm{x}_{\text{cont }}$:
$\mathrm{COH}\left(\mathrm{x}_{\text{cont }},\mathrm{x}_{\text{prompt }}\right)=\frac{\operatorname{EMB}\left(\mathrm{x}_{\text{prompt }}\right)\cdot\operatorname{EMB}\left(\mathrm{x}_{\text{cont }}\right)}{\left\|\operatorname{EMB}\left(\mathrm{x}_{\text{prompt }}\right)\right\|\cdot\left\|\operatorname{EMB}\left(\mathrm{x}_{\text{cont }}\right)\right\|},$
where Emb$(x)$ is the pre-trained SimCSE sentence embedding Gao et al. (2022).
@@ 4.2 Datasets
Following previous studies, we evaluate our proposed method on three domains for open-ended text generation: news, Wikipedia, and stories. For the news domain, we use articles from Wikinews (2000 examples); for the Wikipedia domain, we use the WikiText-103 dataset (1314 examples; Merity et al., 2016); and for the story domain, we use the BookCorpus (Project Gutenberg split, 1947 examples; Zhu et al., 2015). Each of the examples contains a prompt and a gold reference i.e. human-generated continuation for evaluation purposes.
We extract the prompts and decode 256 tokens for the continuations. Finally, we evaluate the generated text based on both the set of metrics (as described in Sec. 4.1) and via human evaluation.
@@ 4.3 Baselines
We compare ACS to widely used decoding methods (including deterministic and stochastic approaches): greedy search, beam search, top-$k$ sampling, nucleus sampling, typical decoding, CD, and CS with constant $\alpha=0.6$ and two versions of $k$: 5 and 10. We include the latter for a fair comparison: The CS generations with $k=10$ achieve better automatic evaluation scores than those generated with $k=5$. Furthermore, they are more comparable to our method, centered around $k=10$. Further, we include an additional adaptive method: DoubleExp, which consists on an exponentiation of the argument of the sigmoid function, with the purpose of reaching values of $\alpha$ closer to 0 or 1. Our goal with this method is to exemplify discrepancies between human judgment and MAUVE. The respective human evaluation results are visualized in Table 2.
@@ 4.4 Models
Table 3: Case Study: Beam search produces degenerative repetitions while nucleus sampling produces text with incoherent semantics w.r.t. the prefix. Contrastive methods produce reasonable repetitions and generate fluent text.
@@ 5.2 Human evaluation
The human evaluation scores for the CS and ACS methods, as well as for CS and DoubleExp, are displayed in Table 2. It is worth mentioning that high MAUVE scores and balanced coherence and diversity do not always align with human judgments. For example, evaluators consistently show a preference for CS over DoubleExp across all datasets. Conversely, ACS is favored in terms of fluency, and AC is preferred for coherence. Nonetheless, when considering all datasets together, there is a slight overall preference for ACS compared to its non-adaptive version.
@@ 5.3 Qualitative examples
We include qualitative examples to highlight the qualitative traits observed for different decoding strategies, exemplified by the generations displayed in Table 3 and the visualizations of the entropy, $k$, and $\alpha$ in Figure 2.
@@ 5.4 Ablation studies
To estimate the effect of different levels of adaptiveness, guided by temperature, on the automatic evaluation we conduct experiments for $q\in\{1,2,4,8,15,20\}$. The results are presented in Table 4, indicating some sensitivity of ACS concerning the choice of $q$, albeit of differing strengths for the different datasets.
@@ 5.5 Generation speed
We have measured the average generation speed across all three datasets by varying the temperature $q\in\{1,2,8\}$ with respect to our baseline (CS with $k=10$ and $\alpha=0.6$). A summary is presented in Table 5, showing a mild decrease in the speed for ACS compared to CS with fixed $\alpha$.
@@
Table 6: Comparion across different languages. Positive $\Delta$-values indicate better performance of ACS vs CS.
@@ 5.6 Application to other languages
We explore the quality of our approach in eight other languages: Arabic, Bengali, German, French, Hindi, Japanese, Dutch, and Chinese, using pre-trained GPT-2-based architectures of different sizes, comparing the performance of ACS to CS. The results are presented in Table 6.
5.7 Effect to varying model sizes
We assess the impact of different model sizes on the quality of text generation, focusing on three model sizes: gpt2-xl, gpt2-large, and gpt2-medium. The generated stories are evaluated across three datasets: Wikinews, Wikitext, and Story. The results, detailed in Table 7, reveal significant discrepancies, particularly with gpt2-medium, where the diversity of CS is notably compromised across all datasets. We suspect this is due to the isotropy of the model, where only high values of $\alpha$ might enable diverse generations. Visual inspection confirms that the outputs are repetitive and degenerate, akin to those produced by greedy search.
@@ 5.8 Findings about MAUVE
Our human evaluation results, as detailed in Table 2, corroborate the findings of Su and Xu (2022). The discrepancies observed between MAUVE scores and human evaluation scores indicate that MAUVE does not consistently align with human preferences. Specifically, the DoubleExp method yields higher MAUVE values compared to CS and ACS. However, human evaluators consistently rate DoubleExp as less coherent and fluent than the CS approach with $k=10$ and $\alpha=0.6$. Moreover, we identified two significant issues related to varying the truncation length in pairwise sentence comparisons: (1) Inconsistent results that lead to different conclusions regarding the optimal decoding method, and (2) substantial differences in sample sizes, as illustrated in Appendix D, Table 8.
@@ 6 Discussion and Future Work
In this study, we introduce a novel approach grounded in a CS framework. It is important to note, however, that the quality of generated text extends beyond just probability and diversity. Other factors, like informativeness, trustworthiness, coherence, and cohesion (as discussed by De Beaugrande and Dressler (1981)), contribute to the overall quality of the text. While some metrics have been proposed previously to measure these attributes (referenced as Work on metrics et al.), there remains potential to integrate them into either the training process or the decoding strategy. Moving forward, our aim is to expand our adaptable method to encompass these traits, thereby improving the evaluation of text quality through a more holistic approach. Furthermore, we would like to explore other criteria for the automatic selection of $k$ and $\alpha$. A path worth exploring could be related to the ratio between generation perplexity and model perplexity, where the algorithm would penalize deviations from the model perplexity through automatic adaptation. Finally, our study has centered on a specific task: Open-ended Text Generation. However, the potential influence of this adaptive decoding strategy on various tasks and contexts warrants further investigation. We aim to broaden our research to evaluate its efficacy in Machine Translation, Summarization, and Handwritten Text Recognition, particularly within the scope of low-resource languages, where our approach may prove beneficial in scenarios where training examples are scarce.
@@ 7 Conclusion
We introduced an uncertainty-guided adaptive method aimed at enhancing the quality of open-ended text generation output. Our approach entails calculating the estimated uncertainty at each time step, assessed through Shannon entropy, and leveraging it to regulate the weighting assigned to model confidence and degeneration penalty, as proposed by Su et al. (2022) and Su and Xu (2022). Our experiments have demonstrated that this method performs well in terms of coherence and diversity, with MAUVE results comparable to existing methods. It receives high ratings from human evaluators, especially for the fluency of its outputs. This method does not require hyperparameter tuning and uses resources similar to the non-adaptive CS. Notably, unlike CD, it does not rely on two separate models, making it more versatile and suitable for various tasks. Furthermore, our comprehensive studies comparing this approach to CS highlight its robustness in multilingualism, speed, and model sizes, albeit with some latency. We hope to see further research in adaptive methods for open-ended text generation and the development of new metrics that better align with human judgments to improve the evaluation of decoding strategies.
@@
The proposed method of ACS could potentially be a step forward in enhancing the quality of text generation from language models. However, it is crucial to acknowledge its limitations and areas for further exploration to ensure a more comprehensive approach to text generation. One primary limitation lies in the narrow focus on only two objectives: model confidence and divergence penalty. While these are certainly critical factors in evaluating text quality, they do not encapsulate the full spectrum of desirable attributes in generated text. Informativeness, fluency, accuracy, trustworthiness, coherence, and cohesion are equally important indicators of high-quality text. Ignoring these aspects in the decoding process could result in text that lacks depth or fails to effectively convey information, particularly in specialized or nuanced contexts.
Another limitation worth considering is the trade-off between speed and quality. While ACS might generate higher-quality text, it does so at the expense of speed. This slower generation process may not be suitable for all applications, especially those requiring real-time responses or large-scale text generation tasks. Balancing the need for quality with the practical constraints of time and computational resources is an ongoing challenge in developing decoding methods for language models.
Additionally, the application of ACS has primarily been explored in the context of text generation tasks such as language modeling and natural language understanding. Its effectiveness on other tasks, such as machine translation, summarization, or multi-modal tasks like handwritten text recognition, remains largely untested. Each of these tasks poses unique challenges and requirements in terms of text diversity, context preservation, and semantic accuracy. Investigating the adaptability and performance of the approach across a broader range of tasks will be essential for its broader applicability and generalizability.
Lastly, while various measures of local uncertainty have been explored in the context of ACS, such as KL divergence, variance, perplexity, and entropy, there is still much room for exploration and refinement. Identifying the most suitable measures and functions for assessing text quality and guiding the decoding process is an ongoing research endeavor. Further experimentation and analysis are needed to determine the optimal combination of uncertainty metrics that can reliably produce high-quality, open-ended text generation across diverse domains and languages. However, the modularity of ACS in terms of choosing the $\delta$-functions allows for a seamless further exploration of different approaches.

### v2

@@ Abstract
Despite the remarkable capabilities of large language models, generating high-quality text remains a challenging task. Numerous decoding strategies—such as beam search, sampling with temperature, top-$k$ sampling, nucleus (top-$p$) sampling, typical decoding, contrastive decoding, and contrastive search—have been proposed to address these challenges by improving coherence, diversity, and resemblance to human-generated text. In this study, we introduce Adaptive Contrastive Search (ACS), a novel decoding strategy that extends contrastive search (CS) by incorporating an adaptive degeneration penalty informed by the model’s estimated uncertainty at each generation step. ACS aims to enhance creativity and diversity while maintaining coherence to produce high-quality outputs. Extensive experiments across various model architectures, languages, and datasets demonstrate that our approach improves both creativity and coherence, underscoring its effectiveness in text-generation tasks. We release our code, datasets, and models to facilitate further research.
@@ 1 Introduction
The Transformer (Vaswani et al., 2017) plays a key role in various generative natural language processing (NLP) tasks, such as generating stories, completing contextual text, and dialogue systems. However, the conventional method of training these models using maximum likelihood estimation (MLE) and decoding to the most probable sequence often results in substantial shortcomings. This can lead to repetitive and uncreative outputs, also known as degenerate text.
Figure 1: Visualization of the Adaptive Contrastive Search (ACS) process: A three-step procedure that uses entropy as a proxy for model uncertainty to automatically adjust contrastive search parameters.
@@
We address this limitation by proposing adaptive contrastive search (ACS), an adaptive approach that automatically adjusts the hyperparameters of conventional CS. This method evaluates the model’s uncertainty at each generation step and adjusts the weighting of both components without the need for manual intervention. The experimental outcomes demonstrate the effectiveness of our method, as it performs well in the task of open-ended text generation across different architectures, languages, and datasets.
@@ 2 Related work
These approaches focus on choosing the text continuation with the highest probability according to the model’s probability distribution. Prominent examples are beam search and greedy search. Recently, studies by Shao et al. (2017); Vijayakumar et al. (2018); Paulus et al. (2017); Klein et al. (2017) have demonstrated that solely maximizing the output probability frequently leads to degenerated or repetitive text sequences, a problem that has been addressed by stochastic, sampling-based methods.
@@
Our study, however, focuses on the work of Su et al. (2022), where they introduce Contrastive Search. In CS, given the prompt text $\boldsymbol{x}_{<t}$, the selection of the output token $x_{t}$ follows:
@@
Empirical results suggest different values for $\alpha$ and $k$, depending on the task, the datasets, and the language of interest, respectively (Su et al., 2022; Su and Xu, 2022; Su and Collier, 2023). An empirical study comparing CS and CD Su and Collier (2023) highlights the strengths and weaknesses of both approaches. The automatic evaluation results suggest that CD performs better on MAUVE Pillutla et al. (2021), while CS excels in diversity and coherence. Additionally, through extensive human evaluations, they demonstrate that human annotators universally prefer CS over CD by substantial margins. Given the contradictory results between MAUVE and human evaluations, their analysis reveals that balancing diversity and coherence metrics better correlates with human judgments.
@@ 3.2 Theoretical Motivation
Table 1: Automatic evaluation results: Numbers marked with * are obtained using the generated texts originally released by Li et al. (2023), CS results are taken from Su and Xu (2022). The highest scores are highlighted in bold.
@@ 4.1 Evaluation Metrics
Proposed by Su et al. (2022), the coherence metric is defined as the averaged log-likelihood of the generated text conditioned on the prompt as
$\operatorname{COH}(\hat{\boldsymbol{x}},\boldsymbol{x})=\frac{1}{|\hat{\boldsymbol{x}}|}\sum_{i=1}^{|\hat{\boldsymbol{x}}|}\log p_{\mathcal{M}}\left(\hat{\boldsymbol{x}}_{i}\mid\left[\boldsymbol{x}:\hat{\boldsymbol{x}}_{<i}\right]\right)$
where $\boldsymbol{x}$ and $\hat{\boldsymbol{x}}$ are the prompt and the generated text, respectively; [:] is the concatenation operation and $\mathcal{M}$ is the OPT model (2.7B) Zhang et al. (2022).
@@ 4.2 Datasets
Following previous studies, we evaluate our proposed method on three domains for open-ended text generation: news, Wikipedia articles, and stories. For the news domain, we use articles from Wikinews (2000 examples); for the Wikipedia domain, we use the WikiText-103 dataset (1314 examples; Merity et al., 2016); and for the story domain, we use the BookCorpus (Project Gutenberg split, 1947 examples; Zhu et al., 2015). Each of the examples contains a prompt and a gold reference i.e. human-generated continuation for evaluation purposes. We extract the prompts and decode 256 tokens for the continuations. Finally, we evaluate the generated text based on both the set of metrics (as described in Sec. 4.1) and human preferences.
@@ 4.3 Baselines
We compare ACS to widely used decoding methods (including deterministic and stochastic approaches): greedy search, beam search, top-$k$ sampling, nucleus sampling, typical decoding, CD, and CS with constant $\alpha=0.6$ and two versions of $k$: 5 and 10. We include the latter for a fair comparison: The CS generations with $k=10$ achieve better automatic evaluation scores than those generated with $k=5$. Furthermore, they are more comparable to our method, centered around $k=10$. Further, we include an additional adaptive method: DoubleExp, which consists on an exponentiation of the argument of the sigmoid function, with the purpose of reaching values of $\alpha$ closer to 0 or 1. Our goal with this method is to exemplify discrepancies between human judgment and MAUVE. The respective human evaluation results are visualized in Table 2 and its implementation is described in Appendix F.
@@ 4.4 Models
Table 3: Case Study: Beam search produces degenerative repetitions while nucleus sampling produces text with incoherent semantics w.r.t. the prefix. Contrastive methods exhibit coherent and fluent text.
@@ 5.2 Human evaluation
The human evaluation scores for the CS and ACS methods, as well as for CS and DoubleExp, are displayed in Table 2. It is worth mentioning that high MAUVE scores do not always align with human judgments. For instance, evaluators consistently show a preference for CS over DoubleExp across all datasets. Conversely, ACS is favored in terms of fluency, and AC is preferred for coherence. Nonetheless, when considering all datasets together, there is a slight overall preference for ACS compared to its non-adaptive version.
@@ 5.3 Qualitative examples
We present qualitative examples to illustrate the distinct characteristics of different decoding strategies. Table 3 highlights the generated text variations, while Figure 2 visualizes the behavior of key parameters such as entropy, $k$, and $\alpha$.
@@ 5.4 Ablation studies
[…]
@@ 5.5 Generation speed
We have measured the average generation speed across all three datasets by varying the temperature $q\in\{1,2,8\}$ with respect to our baseline (CS with $k=10$ and $\alpha=0.6$). A summary is presented in Table 5, showing a decrease of 35% in the speed for ACS compared to CS with fixed $\alpha$. A supplementary analysis of this for smaller values of $k$ is provided in Appendix G.
@@
Table 6: Comparison across different languages. Positive $\Delta$-values indicate better performance of ACS vs CS.
@@ 5.6 Application to other languages
We evaluate the performance of our approach across eight additional languages: Arabic, Bengali, German, French, Hindi, Japanese, Dutch, and Chinese. For this, we utilize pre-trained GPT-2-based architectures of various sizes, comparing the Adaptive Contrastive Search (ACS) to standard Contrastive Search (CS). The detailed results are provided in Table 6. The scores vary greatly across languages and metrics, with particularly strong results in Bengali, French, Hindi, and Chinese. Nevertheless, a clear trend emerges: ACS, with the exception of German, consistently achieves comparable or superior MAUVE scores while maintaining a balance between coherence and diversity, outperforming its static counterpart.
5.7 Effect of varying model sizes
We examine the influence of model size on the quality of text generation, focusing on three different sizes: gpt2-xl (1.50B), gpt2-large (0.76B), and gpt2-medium (0.35B). The generated outputs are evaluated across the Wikinews, Wikitext, and Story datasets. As shown in Table 7, there are marked differences in performance, particularly with gpt2-medium, where the diversity under CS is substantially diminished across all datasets. We hypothesize that this is linked to the model’s isotropy, where only high values of $\alpha$ can foster diversity in the generations. A visual inspection further supports this observation, as the outputs from gpt2-medium tend to be repetitive and degenerate, resembling those produced by greedy or beam search.
@@ 5.8 Findings about MAUVE
Our human evaluation results, as detailed in Table 2, corroborate the findings of Su and Xu (2022). The discrepancies observed indicate that MAUVE does not consistently align with human preferences. Specifically, the DoubleExp method yields higher MAUVE values compared to CS and ACS. However, human evaluators consistently rate DoubleExp as less coherent and fluent than the CS approach with $k=10$ and $\alpha=0.6$. Moreover, we identified two considerable issues related to varying the truncation length in pairwise sentence comparisons: (1) Inconsistent results that lead to different conclusions regarding the optimal decoding method, and (2) substantial differences in sample sizes, as illustrated in Appendix E, Table 8.
5.9 Interpretability
We conducted additional experiments employing CS with varying values of $k$ and $\alpha$, measuring the same automatic metrics for human-generated text and examining which combinations of hyperparameters align most closely with the gold references. In Appendix D, our analysis reveals that extreme parameter settings, such as very low values of $\alpha$, yield very low diversity and MAUVE scores but excessively high coherence (even in combination with high values of $k$). Conversely, very high values of $\alpha$ lead to texts that are overly diverse and incoherent (even with moderate values of $k$). A desired balance emerges from these observations: moderate values of $k$ and $\alpha$, such as $k=10$ and $\alpha=0.6$, tend to approximate human references by favoring both diversity and coherence. Our experiments show that ACS frequently generates results within this range, at times favoring either higher diversity or greater coherence, as dictated by the uncertainty-guided regularization.
@@ 6 Discussion and Future Work
In this study, we introduce a novel approach grounded in a CS framework. It is important to note, however, that the quality of generated text extends beyond just model confidence and diversity. Other factors, like informativeness, trustworthiness, and cohesion (as discussed by De Beaugrande and Dressler (1981)), contribute to the overall quality of the text. Moving forward, our aim is to expand our adaptable method to encompass these traits, thereby improving the evaluation of text quality through a more holistic approach. Furthermore, we would like to explore other criteria for the automatic selection of $k$ and $\alpha$. A path worth exploring could be related to the ratio between generation perplexity and model perplexity, where the algorithm would penalize deviations from the model perplexity through automatic adaptation. Finally, our study has centered on a specific task: Open-ended text generation. However, the potential influence of this adaptive decoding strategy on various tasks and contexts warrants further investigation. We aim to broaden our research to evaluate its efficacy in Machine Translation and Summarization, particularly within the scope of low-resource languages, where our approach may prove beneficial in scenarios where training examples are scarce. Finally, we wish to explore the potential of our approach beyond base models and analyze the effect after SFT and strategic prompting.
@@ 7 Conclusion
We introduce an uncertainty-guided adaptive method aimed at enhancing the quality of open-ended text generation outputs. Our approach calculates the estimated uncertainty at each time step using Shannon entropy and leverages this measure to dynamically adjust the weighting between model confidence and the degeneration penalty, as proposed by Su et al. (2022) and Su and Xu (2022). Our experiments demonstrate that this method performs well in terms of coherence and diversity, achieving MAUVE scores comparable to existing methods. It receives high ratings from human evaluators, particularly for the fluency of its outputs. This method requires no hyperparameter tuning and utilizes computational resources similar to the non-adaptive CS. Notably, unlike CD, it does not rely on two separate models, making it more versatile and suitable for various tasks. While it introduces some latency, our comprehensive studies comparing this approach to CS highlight its robustness across multiple languages and model sizes. We encourage further research into adaptive methods for open-ended text generation and advocate for the development of new metrics that better align with human judgments to improve the evaluation of decoding strategies.
@@
The proposed approach represents a potential step forward in enhancing text generation quality from language models. However, a key limitation lies in the method’s narrow focus on two main objectives: model confidence and degeneration penalty. While these are critical for evaluating certain aspects of text quality, they do not fully capture the broad spectrum of desirable traits in generated text, such as informativeness, fluency, accuracy, trustworthiness, coherence, and cohesion. By not incorporating these additional factors into the decoding process, ACS may produce text that, while coherent and diverse, lacks depth or fails to effectively communicate complex or specialized information (and facts). Expanding the focus to include these dimensions could significantly improve the robustness and applicability of the method across diverse text generation tasks. Another limitation worth considering is the architectural choice of our experiments, which focuses on the family of gpt2-models, in particular in its gpt2-xl version. We plan to extend this analysis to more modern architectures, such as Mistral 7B, (Jiang et al., 2023, 2024), Llama2 7B (Touvron et al., 2023), Llama 3.1 8B (Dubey et al., 2024), Deepseek 7B (DeepSeek-AI et al., 2024), Qwen2 (Yang et al., 2024), Falcon2 (Malartic et al., 2024).
Additionally, the application of ACS has primarily been explored in the context of text generation tasks such as language modeling and natural language understanding. Its effectiveness on other tasks, such as machine translation, summarization, or multi-modal tasks, remains largely untested. Each of these tasks poses unique challenges and requirements in terms of text diversity, context preservation, and semantic accuracy. Investigating the adaptability and performance of the approach across a broader range of tasks will be essential for its broader generalizability. Lastly, while various measures of local uncertainty have been explored in the context of ACS, such as KL divergence, variance, perplexity, and entropy, there is still much room for exploration and refinement. Further experimentation and analysis are needed to determine the optimal combination of uncertainty metrics that can reliably produce high-quality, open-ended text generation across diverse domains and languages. However, the modularity of ACS in terms of choosing the $\delta$-functions allows for a seamless further exploration of different approaches and applications.